Q3. Find the following limits: x1 x+1 (b) lim (c) lim X-3 X-2
Answer #1

(a) lim 7 Lt = x1 (a) lim x²-1 = lim (L-1)(x+1) = (+1) lim (32-1) = 1-1=0 1+1 2 +1 X 1 (b) 8 x ...a363 1 lim 203-23 - 2 - 2 (